# Renewal theory for Brownian motion with stochastically gated targets

### v1

@@ 2.2 Renewal equation
$\rho^{(\epsilon)}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p(x,t|x_{0})+\int_{0}^{t}\rho^{(\epsilon)}_{k|-1}(x,t-\tau|\epsilon)\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)d\tau,$ (2.17)
@@
$\frac{\partial p(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p(x,t|x_{0})}{\partial x^{2}},\quad 0<x<\infty,$ (2.18a)
$p(0,t|x_{0})$ $=0,\quad p(x,0|x_{0})=\delta(x-x_{0}).$ (2.18b)
and $f(x_{0},t)=D\partial_{x}p(0,t|x_{0})$ the associated FPT density for a totally absorbing boundary. The first term on the right-hand side of equation (2.17) represents all sample trajectories that have never hit the boundary at $x=0$ up to time $t$. This is independent of the state of the gate. The corresponding integral represents all trajectories that first reach the boundary at time $\tau$ with the gate closed, which occurs with probability $\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)d\tau$, after which the particle instantaneously resets to position $x=\epsilon$. The particle can then return an arbitrary number of times before encountering an open gate and being absorbed. Indeed, iterating equation (2.17) leads to the series solution
$\rho^{(\epsilon)}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p(x,t|x_{0})$ (2.19)
$+\int_{0}^{t}d\tau\,\Pi_{k,-1}(t-\tau)p(x,t-\tau|\epsilon)\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)$
$+\int_{0}^{t}d\tau\int_{0}^{t-\tau}d\tau^{\prime}\bigg[\Pi_{k,-1}(t-\tau-\tau^{\prime})p(x,t-\tau-\tau^{\prime}|\epsilon)\Pi_{-1,-1}(\tau+\tau^{\prime})$
$\hskip 85.35826pt\times f(\epsilon,\tau^{\prime})\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)\bigg]+\ldots,$
@@
Figure 3: Example sample paths arising in the first three terms of the iterative expansion of the renewal equation (2.19). (The time axis in successive diagrams is rescaled for illustrative purposes.) Each black (red) circle represents hitting the boundary when the gate is closed (open). If the gate is closed then the reflected particle is infinitesimally displaced by an amount $\epsilon$ (see insert).
@@
$\sum_{k=\pm 1}p_{k|k_{0}}(x,t|x_{0})=p(x,t|x_{0}),$ (2.21)
since $\sum_{k=\pm 1}\Pi_{kk_{0}}(t)=1$. Laplace transforming the renewal equation (2.17) with respect to time $t$ and using the convolution theorem then gives
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}_{k|k_{0}}(x,s|x_{0})+\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{f}_{k_{0}}(x_{0},s),$ (2.22)
@@
$\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ $=\int_{0}^{\infty}{\mathrm{e}}^{-st}\Pi_{kk_{0}}(t)p(x,t|x_{0})dt$
@@
$=\sigma_{k}\widetilde{p}(x,s|x_{0})+kk_{0}\sigma_{-k_{0}}\widetilde{p}(x,s+\alpha+\beta|x_{0}).$ (2.23)
@@
$\widetilde{f}_{k_{0}}(x_{0},s)$ $=\sigma_{-1}\widetilde{f}(x_{0},s)-k_{0}\sigma_{-k_{0}}\widetilde{f}(x_{0},s+2\alpha).$ (2.24)
Laplace transforming equations (3.4) we see that $\widetilde{p}(x,s|x_{0})=\widetilde{G}(x,s|x_{0})$, where $G(x,s|x_{0})$ is the Dirichlet Green’s function (2.10), and
$\widetilde{f}(x_{0},s)$ $={D}\partial_{x}\widetilde{G}(0,s|x_{0})={\mathrm{e}}^{-\sqrt{s/D}x_{0}}.$ (2.25)
Next, setting $x_{0}=\epsilon$, $k_{0}=-1$ in equation (2.22) and rearranging shows that
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)=\frac{\widetilde{p}_{k|-1}(x,s|\epsilon)}{1-\widetilde{f}_{-1}(\epsilon,s)}.$ (2.26)
Substituting back into equations (2.22) yields the explicit solution
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}_{k|k_{0}}(x,s|x_{0})+\widetilde{f}_{k_{0}}(x_{0},s)\frac{\widetilde{p}_{k|-1}(x,s|\epsilon)}{1-\widetilde{f}_{-1}(\epsilon,s)}.$ (2.27)
@@
$\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})\equiv\lim_{\epsilon\rightarrow 0}\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}_{k|k_{0}}(x,s|x_{0})+\widetilde{f}_{k_{0}}(x_{0},s)\lim_{\epsilon\rightarrow 0}\frac{\widetilde{p}_{k|-1}(x,s|\epsilon).}{1-\widetilde{f}_{-1}(\epsilon,s)}.$ (2.28)
Since $\widetilde{p}_{k|-1}(x,s|\epsilon)\rightarrow 0$ and $\widetilde{f}_{-1}(\epsilon,s)\rightarrow 1$ as $\epsilon\rightarrow 0$, the limit on the right-hand side of equation (2.28) has to be evaluated using L’hôpital’s rule.
$\lim_{\epsilon\rightarrow 0}\frac{\widetilde{p}_{k|-1}(x,s|\epsilon)}{1-\widetilde{f}_{-1}(\epsilon,s)}=\frac{\sigma_{k}{\mathrm{e}}^{-\sqrt{s/D}x}-k\sigma_{1}{\mathrm{e}}^{-\sqrt{s+2\alpha]/D}x}}{\sigma_{-1}\sqrt{sD}+\sigma_{1}\sqrt{[s+2\alpha]D}}.$ (2.29)
It can be checked that the solution (2.28) for $k=\pm 1$ is equivalent to equation (2.1), which is the solution obtained directly from the forward Kolmogorov equation.
We also note that summing both sides of equation (2.28) with respect to $k=\pm 1$ and integrating with respect to $x\in[0,\infty)$ gives the following equation for the stochastically gated survival probability in Laplace space:
$\widetilde{\mathfrak{Q}}_{k_{0}}(x_{0},s)=\widetilde{Q}(x_{0},s)+\widetilde{f}_{k_{0}}(x_{0},s)\lim_{\epsilon\rightarrow 0}\frac{\widetilde{Q}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)},$ (2.30)
where (see equation (2.5))
$\widetilde{{\mathfrak{Q}}}_{k_{0}}(x_{0},s)$ $:=\int_{0}^{\infty}\sum_{k=\pm 1}\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})dx$ (2.31)
@@
$\widetilde{Q}(\epsilon,s)=\sum_{k=\pm 1}\int_{0}^{\infty}\widetilde{p}_{k|-1}(x,s|\epsilon)dx$ $=\int_{0}^{\infty}p(x,s|\epsilon)dx$ (2.32)
is the Laplace transformed survival probability with respect to adsorption. We have reversed the order of integration and the limit $\epsilon\rightarrow 0$ in equation (2.30). Again the limit is evaluated using L’Hôpital’s rule. In the case of the half-line, we have $\widetilde{Q}=\widetilde{Q}_{0}$ with
$\widetilde{Q}_{0}(x_{0},s)$ $:=\int_{0}^{\infty}\widetilde{G}(x,s|x_{0})dx=\frac{1-{\mathrm{e}}^{-\sqrt{s/D}x_{0}}}{s}$ (2.33)
@@
$\lim_{\epsilon\rightarrow 0}\ \frac{\widetilde{Q}_{0}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)}.=\frac{1}{\sigma_{-1}s+\sigma_{1}\sqrt{[s+\alpha+\beta]s}}.$ (2.34)
@@
(i) The renewal equation (2.17) provides a general probabilistic formulation of a stochastically gated boundary. A major advantage compared to the forward Kolmogorov equation (2.4a-c), is that the renewal equation explicitly separates the FPT problem for detecting the gated boundary from the mechanism for subsequently restarting the diffusion process. For example, the particle could reset to its initial position $x_{0}$ whenever it reflects off a closed gate, analogous to desorption-induced resetting [46]. This means setting $\epsilon=x_{0}$ in equation (2.30) rather than taking the limit $\epsilon\rightarrow 0$. In addition, the basic structure of the renewal equation still holds under various modifications of the diffusion process prior to each adsorption event. The only difference is in the construction of the densities $p(x,t|x_{0})$ and $f(x_{0},t)$. For example, the particle could reset to its initial position at a random sequence of times generated by a Poisson process (see section 3). This bulk resetting is distinct from desorption-induced resetting. Alternatively, the condition for terminating each round of bulk diffusion could be modified by setting $p=p^{(\kappa_{0})}$ where $p^{(\kappa_{0})}$ satisfies a Robin boundary condition with reactivity $\kappa_{0}$:
$\frac{\partial p^{(\kappa_{0})}(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p^{(\kappa_{0})}((x,t|x_{0})}{\partial x^{2}},\quad 0<x<\infty,$ (2.35a)
$D\frac{\partial p^{(\kappa_{0})}((0,t|x_{0})}{\partial x}$ $=\kappa_{0}p^{(\kappa_{0})}(0,t|x_{0}),\quad p^{(\kappa_{0})}(x,0|x_{0})=\delta(x-x_{0}).$ (2.35b)
In contrast to the case of a totally adsorbing boundary ($\kappa_{0}\rightarrow\infty$), if the particle is adsorbed when the gate is closed then the diffusive process can be restarted at $x=0$ without any shift in position. The renewal equation (2.17) thus becomes
$\rho^{(\kappa_{0})}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p^{(\kappa_{0})}(x,t|x_{0})$
$+\int_{0}^{t}\rho^{(\kappa_{0})}_{k|-1}(x,t-\tau|0)\Pi_{-1,k_{0}}(\tau)f^{(\kappa_{0})}(x_{0},\tau)d\tau,$ (2.36)
@@
$\widetilde{\rho}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})+\widetilde{f}^{(\kappa_{0})}_{k_{0}}(x_{0},s)\frac{\widetilde{p}_{k|-1}^{(\kappa_{0})}(x,s|0)}{1-\widetilde{f}^{(\kappa_{0})}_{-1}(0,s)}.$ (2.37)
@@
Figure 4: Schematic illustration comparing (a) partial absorption through an open gate and (b) partial adsorption prior to encountering the gate. Reduction of a thick arrow indicates partial reflection.
(iii) The renewal equation (2.36) for finite $\kappa_{0}$ assumes that there is a necessary intermediate step (partial adsorption) prior to detecting the state of the gate, see Fig. 4(a). This differs from the classical treatment of stochastically gated chemical reactions developed by Szabo et al. [6, 14], see also Ref. [24]. These authors treated the closed gate as totally reflecting and the open gate as partially absorbing, see Fig. 4(b). In other words, $\rho^{(\kappa_{0})}_{-1|k_{0}}(x,t|x_{0})$ and $\rho^{(\kappa_{0})}_{1|k_{0}}(x,t|x_{0})$ satisfiied Neumann and Robin boundary conditions, respectively. Green’s function methods were then used to construct an integral equation relating the propagator for $\kappa_{0}>0$ to the propagator for a totally reflecting boundary ($\kappa_{0}=0$). It is instructive to determine the corresponding boundary conditions satisfied by the solution of the renewal equation (2.37) when $\kappa_{0}$ is finite. First note that $\widetilde{f}^{(\kappa_{0})}_{k_{0}}(x_{0},s)=\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(0,s|x_{0})$. Second, the solution of equations (2.35) is [10]
$\widetilde{p}(x,s|x_{0})=\frac{1}{2\sqrt{sD}}\left({\mathrm{e}}^{-\sqrt{s/D}|x-x_{0}|}+\frac{\sqrt{sD}-\kappa_{0}}{\sqrt{sD}+\kappa_{0}}{\mathrm{e}}^{-\sqrt{s/D}(x+x_{0})}\right),$ (2.38)
@@
$\widetilde{p}(0,s|x_{0})=\frac{1}{\sqrt{sD}+\kappa_{0}}{\mathrm{e}}^{-\sqrt{s/D}x_{0}}.$ (2.39)
@@
$\left.D\frac{\partial\widetilde{p}(x,s|0)}{\partial x}\right|_{x=0}=\kappa_{0}\widetilde{p}(0,s|0)-1.$ (2.40)
@@
$\left.D\frac{\partial\widetilde{\rho}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})}{\partial x}\right|_{x=0}=\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{k|k_{0}}(0,s|x_{0})+\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(0,s|x_{0})\frac{\kappa_{0}\widetilde{p}_{k|-1}^{(\kappa_{0})}(0,s|0)-1}{1-\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{-1|-1}(0,s|0)}.$ (2.41)
@@
$\left.D\frac{\partial\widetilde{\rho}^{(\kappa_{0})}_{-1|k_{0}}(x,s|x_{0})}{\partial x}\right|_{x=0}=0,$ (2.42)
@@
$\alpha={(1-\eta)\gamma_{0}}/{\delta},\quad\beta={\eta\gamma_{0}}/{\delta},$ (2.43)
with $0<\eta<1$ and $0<\delta\ll 1$. Ignoring short-time transients, we can take the Laplace variable $s$ to be bounded from above and $\delta$ to be sufficiently small so that $\alpha+\beta\gg s$. From equations (2.9) and (2.15) we have
$q(0,s)=A(s),\quad D\partial_{x}q(0,s)={\mathrm{e}}^{-\sqrt{s/D}x_{0}}-\sqrt{sD}A(s)$ (2.44)
@@
$A(s)\sim\frac{(1-\eta){\mathrm{e}}^{-\sqrt{s/D}x_{0}}}{\eta\sqrt{(\alpha+\beta)D}}=\sqrt{\frac{\delta}{\gamma_{0}D}}\frac{(1-\eta){\mathrm{e}}^{-\sqrt{s/D}x_{0}}}{\eta}$ (2.45)
@@
$D\partial_{x}q(0,s)\sim{\mathrm{e}}^{-\sqrt{s/D}x_{0}}\sim\sqrt{\frac{\gamma_{0}D}{\delta}}\frac{\eta}{1-\eta}q(0,s)$ (2.46)
@@
$\eta\sim\kappa_{\rm eff}\sqrt{{\delta}/{\gamma_{0}D}}.$ (2.47)

### v2

@@ 2.2 Renewal equation
$\rho^{(\epsilon)}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p(x,t|x_{0})+\int_{0}^{t}\rho^{(\epsilon)}_{k|-1}(x,t-\tau|\epsilon)\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)d\tau,$ (2.4gr)
@@
$\frac{\partial p(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p(x,t|x_{0})}{\partial x^{2}},\quad 0<x<\infty,$ (2.4gsa)
$p(0,t|x_{0})$ $=0,\quad p(x,0|x_{0})=\delta(x-x_{0}),$ (2.4gsb)
and $f(x_{0},t)=D\partial_{x}p(0,t|x_{0})$ the associated FPT density for a totally absorbing boundary. The first term on the right-hand side of equation (2.4gr) represents all sample trajectories that have never hit the boundary at $x=0$ up to time $t$. This is independent of the state of the gate. The corresponding integral represents all trajectories that first reach the boundary at time $\tau$ with the gate closed, which occurs with probability $\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)d\tau$, after which the particle instantaneously resets to position $x=\epsilon$. The particle can then return an arbitrary number of times before encountering an open gate and being absorbed. Indeed, iterating equation (2.4gr) leads to the series solution
$\rho^{(\epsilon)}_{k|k_{0}}(x,t|x_{0})$
$=\Pi_{kk_{0}}(t)p(x,t|x_{0})+\int_{0}^{t}d\tau\,\Pi_{k,-1}(t-\tau)p(x,t-\tau|\epsilon)\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)$
$+\int_{0}^{t}d\tau\int_{0}^{t-\tau}d\tau^{\prime}\bigg[\Pi_{k,-1}(t-\tau-\tau^{\prime})p(x,t-\tau-\tau^{\prime}|\epsilon)\Pi_{-1,-1}(\tau+\tau^{\prime})f(\epsilon,\tau^{\prime})$
$\hskip 85.35826pt\times\Pi_{-1,k_{0}}(\tau)f(x_{0},\tau)\bigg]+\ldots,$ (2.4gst)
@@
Figure 3: Example sample paths arising in the first three terms of the iterative expansion of the renewal equation (2.4gst). (The time axis in successive diagrams is rescaled for illustrative purposes.) Each black (red) circle represents hitting the boundary when the gate is closed (open). If the gate is closed then the reflected particle is infinitesimally displaced by an amount $\epsilon$ (see insert).
@@
$\sum_{k=\pm 1}p_{k|k_{0}}(x,t|x_{0})=p(x,t|x_{0}),$ (2.4gsv)
since $\sum_{k=\pm 1}\Pi_{kk_{0}}(t)=1$. Laplace transforming the renewal equation (2.4gr) with respect to time $t$ and using the convolution theorem then gives
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}_{k|k_{0}}(x,s|x_{0})+\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)\widetilde{f}_{k_{0}}(x_{0},s),$ (2.4gsw)
@@
$\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ $=\int_{0}^{\infty}{\mathrm{e}}^{-st}\Pi_{kk_{0}}(t)p(x,t|x_{0})dt$ (2.4gsx)
@@
$=\sigma_{k}\widetilde{p}(x,s|x_{0})+kk_{0}\sigma_{-k_{0}}\widetilde{p}(x,s+\alpha+\beta|x_{0}).$
@@
$\widetilde{f}_{k_{0}}(x_{0},s)$ $=\sigma_{-1}\widetilde{f}(x_{0},s)-k_{0}\sigma_{-k_{0}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widetilde{f}(x_{0},s+\alpha+\beta)}.$ (2.4gsy)
Laplace transforming equations (2.4gsa) and (2.4gsb) we see that $\widetilde{p}(x,s|x_{0})=\widetilde{G}(x,s|x_{0})$, where $G(x,s|x_{0})$ is the Dirichlet Green’s function (2.4gk), and
$\widetilde{f}(x_{0},s)$ $={D}\partial_{x}\widetilde{G}(0,s|x_{0})={\mathrm{e}}^{-\sqrt{s/D}x_{0}}.$ (2.4gsz)
Next, setting $x_{0}=\epsilon$, $k_{0}=-1$ in equation (2.4gsw) and rearranging shows that
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(x,s|\epsilon)=\frac{\widetilde{p}_{k|-1}(x,s|\epsilon)}{1-\widetilde{f}_{-1}(\epsilon,s)}.$ (2.4gsaa)
Substituting back into equations (2.4gsw) yields the explicit solution
$\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}_{k|k_{0}}(x,s|x_{0})+\widetilde{f}_{k_{0}}(x_{0},s)\frac{\widetilde{p}_{k|-1}(x,s|\epsilon)}{1-\widetilde{f}_{-1}(\epsilon,s)}.$ (2.4gsab)
@@
$\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})\equiv\lim_{\epsilon\rightarrow 0}\widetilde{\rho}^{(\epsilon)}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}_{k|k_{0}}(x,s|x_{0})+\widetilde{f}_{k_{0}}(x_{0},s)\lim_{\epsilon\rightarrow 0}\frac{\widetilde{p}_{k|-1}(x,s|\epsilon).}{1-\widetilde{f}_{-1}(\epsilon,s)}.$ (2.4gsac)
Since $\widetilde{p}_{k|-1}(x,s|\epsilon)\rightarrow 0$ and $\widetilde{f}_{-1}(\epsilon,s)\rightarrow 1$ as $\epsilon\rightarrow 0$, the limit on the right-hand side of equation (2.4gsac) has to be evaluated using L’hôpital’s rule.
$\lim_{\epsilon\rightarrow 0}\frac{\widetilde{p}_{k|-1}(x,s|\epsilon)}{1-\widetilde{f}_{-1}(\epsilon,s)}=\frac{\sigma_{k}{\mathrm{e}}^{-\sqrt{s/D}x}-k\sigma_{1}{\mathrm{e}}^{-\sqrt{s+2\alpha]/D}x}}{\sigma_{-1}\sqrt{sD}+\sigma_{1}\sqrt{[s+2\alpha]D}}.$ (2.4gsad)
It can be checked that the solution (2.4gsac) for $k=\pm 1$ is equivalent to equation (2.4gq), which is the solution obtained directly from the forward Kolmogorov equation.
We also note that summing both sides of equation (2.4gsac) with respect to $k=\pm 1$ and integrating with respect to $x\in[0,\infty)$ gives the following equation for the stochastically gated survival probability in Laplace space:
$\widetilde{\mathfrak{Q}}_{k_{0}}(x_{0},s)=\widetilde{Q}(x_{0},s)+\widetilde{f}_{k_{0}}(x_{0},s)\lim_{\epsilon\rightarrow 0}\frac{\widetilde{Q}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)},$ (2.4gsae)
where (see equation (2.4e))
$\widetilde{{\mathfrak{Q}}}_{k_{0}}(x_{0},s)$ $:=\int_{0}^{\infty}\sum_{k=\pm 1}\widetilde{\rho}_{k|k_{0}}(x,s|x_{0})dx,$ (2.4gsaf)
@@
$\widetilde{Q}(\epsilon,s)=\sum_{k=\pm 1}\int_{0}^{\infty}\widetilde{p}_{k|-1}(x,s|\epsilon)dx$ $=\int_{0}^{\infty}p(x,s|\epsilon)dx$ (2.4gsag)
is the Laplace transformed survival probability with respect to adsorption. We have reversed the order of integration and the limit $\epsilon\rightarrow 0$ in equation (2.4gsae). Again the limit is evaluated using L’Hôpital’s rule. In the case of the half-line, we have $\widetilde{Q}=\widetilde{Q}_{0}$ with
$\widetilde{Q}_{0}(x_{0},s)$ $:=\int_{0}^{\infty}\widetilde{G}(x,s|x_{0})dx=\frac{1-{\mathrm{e}}^{-\sqrt{s/D}x_{0}}}{s},$ (2.4gsah)
@@
$\lim_{\epsilon\rightarrow 0}\ \frac{\widetilde{Q}_{0}(\epsilon,s)}{1-\widetilde{f}_{-1}(\epsilon,s)}.=\frac{1}{\sigma_{-1}s+\sigma_{1}\sqrt{[s+\alpha+\beta]s}}.$ (2.4gsai)
@@
(i) The renewal equation (2.4gr) provides a general probabilistic formulation of a stochastically gated boundary. A major advantage compared to the forward Kolmogorov equation (2.4a)–(2.4c), is that the renewal equation explicitly separates the FPT problem for detecting the gated boundary from the mechanism for subsequently restarting the diffusion process. For example, the particle could reset to its initial position $x_{0}$ whenever it reflects off a closed gate, analogous to desorption-induced resetting [46]. This means setting $\epsilon=x_{0}$ in equation (2.4gsae) rather than taking the limit $\epsilon\rightarrow 0$. In addition, the basic structure of the renewal equation still holds under various modifications of the diffusion process prior to each adsorption event. The only difference is in the construction of the densities $p(x,t|x_{0})$ and $f(x_{0},t)$. For example, the particle could reset to its initial position at a random sequence of times generated by a Poisson process (see section 3). This bulk resetting is distinct from desorption-induced resetting. Alternatively, the condition for terminating each round of bulk diffusion could be modified by setting $p=p^{(\kappa_{0})}$ where $p^{(\kappa_{0})}$ satisfies a Robin boundary condition with reactivity $\kappa_{0}$:
$\frac{\partial p^{(\kappa_{0})}(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p^{(\kappa_{0})}((x,t|x_{0})}{\partial x^{2}},\quad 0<x<\infty,$ (2.4gsaja)
$D\frac{\partial p^{(\kappa_{0})}((0,t|x_{0})}{\partial x}$ $=\kappa_{0}p^{(\kappa_{0})}(0,t|x_{0}),\quad p^{(\kappa_{0})}(x,0|x_{0})=\delta(x-x_{0}).$ (2.4gsajb)
In contrast to the case of a totally adsorbing boundary ($\kappa_{0}\rightarrow\infty$), if the particle is adsorbed when the gate is closed then the diffusive process can be restarted at $x=0$ without any shift in position. The renewal equation (2.4gr) thus becomes
$\rho^{(\kappa_{0})}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p^{(\kappa_{0})}(x,t|x_{0})+\int_{0}^{t}\rho^{(\kappa_{0})}_{k|-1}(x,t-\tau|0)\Pi_{-1,k_{0}}(\tau)f^{(\kappa_{0})}(x_{0},\tau)d\tau,$
@@
$\widetilde{\rho}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})+\widetilde{f}^{(\kappa_{0})}_{k_{0}}(x_{0},s)\frac{\widetilde{p}_{k|-1}^{(\kappa_{0})}(x,s|0)}{1-\widetilde{f}^{(\kappa_{0})}_{-1}(0,s)}.$ (2.4gsajal)
@@
Figure 4: Schematic illustration comparing (a) partial adsorption prior to encountering the gate and (b) partial absorption through an open gate. Reduction of a thick arrow indicates partial reflection.
(iii) The renewal equation (LABEL:renkap) for finite $\kappa_{0}$ assumes that there is a necessary intermediate step (partial adsorption) prior to detecting the state of the gate, see Fig. 4(a). This differs from the classical treatment of stochastically gated chemical reactions developed by Szabo et al. [6, 14], see also Ref. [24]. These authors treated the closed gate as totally reflecting and the open gate as partially absorbing, see Fig. 4(b). In other words, $\rho^{(\kappa_{0})}_{-1|k_{0}}(x,t|x_{0})$ and $\rho^{(\kappa_{0})}_{1|k_{0}}(x,t|x_{0})$ satisfiied Neumann and Robin boundary conditions, respectively. Green’s function methods were then used to construct an integral equation relating the propagator for $\kappa_{0}>0$ to the propagator for a totally reflecting boundary ($\kappa_{0}=0$). It is instructive to determine the corresponding boundary conditions satisfied by the solution of the renewal equation (2.4gsajal) when $\kappa_{0}$ is finite. First note that $\widetilde{f}^{(\kappa_{0})}_{k_{0}}(x_{0},s)=\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(0,s|x_{0})$. Second, the solution of equations (2.4gsaja) and (2.4gsajb) is [10]
$\widetilde{p}(x,s|x_{0})=\frac{1}{2\sqrt{sD}}\left({\mathrm{e}}^{-\sqrt{s/D}|x-x_{0}|}+\frac{\sqrt{sD}-\kappa_{0}}{\sqrt{sD}+\kappa_{0}}{\mathrm{e}}^{-\sqrt{s/D}(x+x_{0})}\right),$ (2.4gsajam)
@@
$\widetilde{p}(0,s|x_{0})=\frac{1}{\sqrt{sD}+\kappa_{0}}{\mathrm{e}}^{-\sqrt{s/D}x_{0}}.$ (2.4gsajan)
@@
$\left.D\frac{\partial\widetilde{p}(x,s|0)}{\partial x}\right|_{x=0}=\kappa_{0}\widetilde{p}(0,s|0)-1.$ (2.4gsajao)
@@
$\left.D\frac{\partial\widetilde{\rho}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})}{\partial x}\right|_{x=0}=\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{k|k_{0}}(0,s|x_{0})+\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(0,s|x_{0})\frac{\kappa_{0}\widetilde{p}_{k|-1}^{(\kappa_{0})}(0,s|0)-1}{1-\kappa_{0}\widetilde{p}^{(\kappa_{0})}_{-1|-1}(0,s|0)}.$ (2.4gsajap)
@@
$\left.D\frac{\partial\widetilde{\rho}^{(\kappa_{0})}_{-1|k_{0}}(x,s|x_{0})}{\partial x}\right|_{x=0}=0,$ (2.4gsajaq)
@@
$\alpha={(1-\eta)\gamma_{0}}/{\delta},\quad\beta={\eta\gamma_{0}}/{\delta},$ (2.4gsajar)
with $0<\eta<1$ and $0<\delta\ll 1$. Ignoring short-time transients, we can take the Laplace variable $s$ to be bounded from above and $\delta$ to be sufficiently small so that $\alpha+\beta\gg s$. From equations (2.4gj) and (2.4gp) we have
$q(0,s)=A(s),\quad D\partial_{x}q(0,s)={\mathrm{e}}^{-\sqrt{s/D}x_{0}}-\sqrt{sD}A(s),$ (2.4gsajas)
@@
$A(s)\sim\frac{(1-\eta){\mathrm{e}}^{-\sqrt{s/D}x_{0}}}{\eta\sqrt{(\alpha+\beta)D}}=\sqrt{\frac{\delta}{\gamma_{0}D}}\frac{(1-\eta){\mathrm{e}}^{-\sqrt{s/D}x_{0}}}{\eta}.$ (2.4gsajat)
@@
$D\partial_{x}q(0,s)\sim{\mathrm{e}}^{-\sqrt{s/D}x_{0}}\sim\sqrt{\frac{\gamma_{0}D}{\delta}}\frac{\eta}{1-\eta}q(0,s).$ (2.4gsajau)
@@
$\eta\sim\kappa_{\rm eff}\sqrt{{\delta}/{\gamma_{0}D}}.$ (2.4gsajav)